OpenStudy (anonymous):

simplify x^-2 y^0 over 2^-1 w^0 z^3 A.-2x^3z^3 B. x^2 z^3 over 2 C. 1 over 2x^2 z^3 D. 2 over x^2 z^3 please help?

OpenStudy (anonymous):

hey luigi how ya doin bud can you help?

OpenStudy (luigi0210):

First thing, anything to the 0 is 1

OpenStudy (anonymous):

ik I jus put the problem up there as I saw it bc I didn't kno it

OpenStudy (luigi0210):

you remember earlier how anything negative changes places?

OpenStudy (anonymous):

yes I wanted to say C but didn't know.. :(

OpenStudy (anonymous):

because the -1 was a +1

OpenStudy (luigi0210):

\[\frac{ x^{-2}y^0 }{ 2^{-1}w^0z^3 }\]

OpenStudy (anonymous):

That the question just to simplify that

OpenStudy (anonymous):

Lap top died sorry for the delay

OpenStudy (luigi0210):

Let's start by changing all those 0's to ones, so they basically don't exist leaving: \[\frac{ x^{-2} }{ z^{-3}2^{-1} }\]

OpenStudy (luigi0210):

ignore that negative z

OpenStudy (luigi0210):

\[\frac{ x^{-2} }{ z^32^{-1} }\]

OpenStudy (anonymous):

Change the negative exponents to positive right?

OpenStudy (luigi0210):

Yup, so do it and show me what you get

OpenStudy (anonymous):

X^2 over z^3 2^1

OpenStudy (luigi0210):

Ah, remember that they switch too.. so it was ontop it would be on bottom and visa versa

OpenStudy (anonymous):

Which answer would it be though?

OpenStudy (luigi0210):

It would be D

OpenStudy (anonymous):

Thanks bud

OpenStudy (luigi0210):

No problem